OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. A. 84 B. 7 C. -12

OpenStudy (anonymous):

y = k*x -7 = k (-1) k = 7 y = 7x -84 = 7x x = -84/7 = 12

OpenStudy (anonymous):

-12*

OpenStudy (anonymous):

thanks